Write each of the following as decimals: 
Two tens and nine-tenths

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

The decimal form of two tens and nine tenths is 20 + 9 / 10 = 20.9

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
The decimal form of "Two tens and nine tenths " can be expressed as 20+9/10 = 209/10 = 20.9
 
Hope it helps!
Thank you ✌️